the half life period of 1st order reaction is 10 minutes .the time required for the concentraton of reactant to change from 0.08M to 0.06M

Grade:12

2 Answers

Arun
4 years ago
K = 0.693/t1/2
k = 0.693/10 = 0.0693
 
ln (0.08/0.02) = kt
 
ln 4 = kt
 
hence t = 20 minutes
 
Hope it helps
 
Vikas TU
4 years ago
Dear student 
The above ans is incorrect.
We know that half -life is related to K as
K = 0.693t1/2
given t1 /2 = 10 min
thus K = 0,693 x 10 min = 0.0693min-1
t = 2.303/K log [R0]/[R]
= 2.303/0.0693 log [0.08]/[0.06]
= 4.15 minute.